Urgent PLZZ

Supply Club, Inc., sells a variety of paper products, office supplies, and other products used by businesses and individual consumers. During July 2021 it started a loyalty program through which qualifying customers can accumulate points and redeem those points for discounts on future purchases. Redemption of a loyalty point reduces the price of one dollar of future purchases by 20% (equal to 20 cents). Customers do not earn additional loyalty points for purchases on which loyalty points are redeemed. Based on past experience, Supply Club estimates a 60% probability that any point issued will be redeemed for the discount. During July 2021, the company records $161,500 of revenue and awards 237,500 loyalty points. The aggregate stand-alone selling price of the purchased products is $161,500. Seventy-five percent of sales were cash sales, and the remainder were credit sales.

Required:
1. & 2. Prepare Supply Club’s journal entry to record July and August sales. During August, customers redeem loyalty points on $114,000 of merchandise. Seventy-five percent of those sales were for cash, and the remainder were credit sales. (Do not round intermediate calculations. If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)
Answer #1

Answer:

1 & 2 ) Prepare Supply Club's journal entry to record july and August sales.

Date Particulars Debit ($) Credit ($)
31.07.2021 Cash (161500*0.75 = 121125 121,125
Accounts Receivable 32,300
Revenue 137,275
Deferred Revenue 24,225

Working Notes:
Loyalty points: (0.20*0.60*237500) = 28500
Product sales: 161500
Allocation:
Loyalty points: 28500 / (28500+161500) = 0.15*161500 = 24225
Product sales: 161500 / (28500+161500) = 0.85*161500 = 137275

Date Particulars Debit ($) Credit ($)
31.08.2021 Cash (114000*(1-0.2)*0.75=68400) 68,400
Accounts Receivable (114000*(1-0.2)*0.25=22800 22800
Deferred Revenue 19380
Revenue 110580

Working Notes:
Deferred Revenue:
237500 points issued*60% expected to be redeemed = 142500 points to be redeemed
114000 points redeemed / 142500 expected = 80% redeemed
$24,225 deferred revenue*80% = $19380 Revenue earned